OpenStudy (anonymous):

Fan and Medal Which of the following is not a perfect square trinomial? A. 144 + 12y + y^2 B. 100 + 20y + y^2 C. 64 – 16y + y^2 D. 25 + 10y + y^2

OpenStudy (anonymous):

1: 144 +24y +y^2 = y^2 +24y +144 = (y+12)(y+12) 2: 100 -20y +y^2 = y^2 -20y +100 = (y-10)(y-10) 3: 81 +18y +y^2 = y^2 +18y +81 = (y+9)(y+9) 4: 121 +11y +y^2 = y^2 +11y +121 = *** does not factor **** 5: m^2 +12m +36 = (m+6)(m+6)

OpenStudy (anonymous):

D isn't a perfect square trinomial